OpenStudy (anonymous):

verify that the function f(X)=x^3-7 and g(X)=3sqrt x+7 are inverse of each other

OpenStudy (lgbasallote):

do you know how to solve for f(g(x))?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

good. first, let me tell you the concept they are inverse if: f(g(x)) = g(f(x)) so to verify...you need to solve for f(g(x)) and g(f(x)) and see if they are equal does that help?

OpenStudy (anonymous):

i try first , k?

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

f[g(X)]=f(3sqr [x+7]]

OpenStudy (lgbasallote):

yes go on... substitute \(3\sqrt{x+7}\) into the x of x^3 - 7

OpenStudy (anonymous):

f(g(X))=(3sqrt x+7)^3 -7 , i dont no how to solve this

OpenStudy (lgbasallote):

one question first... is that \[3\sqrt{x+7}\]or \[3\sqrt x + 7\]

OpenStudy (anonymous):

the first one

OpenStudy (lgbasallote):

oh cool. anyway...try solving g(f(x)) first if you're stuck here. you dont know...it might look like that too

sam (.sam.):

Actually you can just solve for f(X)=x^3-7 Let y=x^3-7 y+7=x^3 \[\sqrt[3]{y+7}=x\] \[f^{-1}(x)=\sqrt[3]{x+7}=g(x)\]